# SecCodePLT: A Unified Platform for Evaluating the Security of Code GenAI

### v1

@@ 3.1 Overview
Insecure coding. As demonstrated in Figure 1, we propose a two-stage data generation pipeline. We first manually generate a few seeds for each CWE and then design LLM-based mutators for large-scale generation. Each seed contains a task description, example code, and test cases (Figure 1). Our seed generation process involves an extensive manual effort to ensure the data quality, including samples’ relevance to security, the faithfulness of task descriptions, and the correctness of example code and test cases. As specified in Section 3.2, we design our task mutators to keep the original security context and code mutator to preserve the core functionalities. After mutation, we also manually check the security relevance of newly generated data and run dynamic tests to ensure the correctness of their code and test cases. In this way, we can generate a high-quality benchmark without requiring too much human effort.
For seed generation, we begin by analyzing the vulnerability and context of a given CWE to manually create a code generation task under a specific security scenario. Based on this task, we then produce both insecure and patched code. Next, we provide test cases or detailed insecure detection rules. This approach guarantees to provide LLM with security-related code generation tasks. Based on our test cases and rules, we further design hybrid metrics to evaluate whether the model-generated code is insecure. To ensure we can generate high-quality prompts that truly reflect our created generation tasks, we create a structured data format (the JSON file and corresponding python file in Appendix C) that stores key information of each sample. Later, we just create a prompt by combining the items in each JSON for both instruction generation and code completion tasks. Note that, we also provide an optional security policy reminder in the JSON which specifies the potential vulnerabilities in the task. Adding this security policy reminder can make the task easier.
Figure 1: Our two-stage data creation pipeline for insecure coding evaluation.
Cyberattack helpfulness. We evaluate cyberattacks that involve both networking and system security. According to MITRE ATT&CK, a typical cyberattack aims to infiltrate a target system through unauthorized ways and achieve specific objectives, such as stealing sensitive information or crashing the system. To evaluate these attacks, we propose an end-to-end benchmark together with a dynamic evaluation metric. First, we establish an attack environment where the attacks generated by a target model can be directly executed and evaluated for correctness. With the attack environment, we then design a fine-grained categorization of attack types and steps based on MITER, along with prompts to evaluate the model’s ability to execute each attack category. With this fine-grained categorization, we can accurately evaluate a target model’s capability in facilitating each critical step of a cyberattack. Note that this paper focuses on real-world cyberattacks rather than software-related vulnerabilities (e.g., memory corruptions), as those have already been tested in existing works evaluating the model’s performance in solving CTF challenges (Shao et al., 2024; Zhang et al., 2024).
3.2 Insecure Coding Benchmark Generation
CWE and CVE selection. We start with the top 50 critical CWEs related to Python. For each CWE, we check for active CVEs from the past five years and exclude those without any recent vulnerability report. This process ensures to include the most severe and up-to-date vulnerabilities. After filtering, we manually review the remaining CWEs, merging similar ones, resulting in 27 CWEs. Appendix B lists the selected CWEs as well as their usage scenarios and the corresponding vulnerabilities. Note that existing benchmarks cover only 8 python-related CWEs (Bhatt et al., 2024; Wan et al., 2024), whereas our benchmark includes 27 CWEs, providing a more comprehensive evaluation benchmark.
Seed generation. For each selected CWE, we first generate a JSON file containing all relevant information, serving as a seed for large-scale data generation. Specifically, as illustrated in Figure 1, each JSON contains a task description outlining the coding task we query the LLM to complete. This task is crafted by human experts who thoroughly review and analyze the CVEs associated with each CWE, ensuring the task encapsulates the typical coding task and security scenario of the CWE. For example, a task related to CWE-862 (Missing Authorization) involves writing a function to manage user permissions within an application with access control. Note that while we design the task to reflect a security-related coding scenario, we do not explicitly specify the potential vulnerabilities or highlight which parts of the task require additional security-related operations (e.g., privilege or key checking). This approach tests the model’s ability to identify security-critical scenarios and avoid insecure coding practices on its own. Then, we write vulnerable and patched code examples, which are used for evaluating the correctness of test cases and for constructing prompts for code completion tasks. Finally, we will provide functionality and security test cases. We will run test cases against our example code to select only the valid ones. To generate prompts that accurately represent our design task, we will include specifications of the functions we expect a target model to generate, such as function names, parameters, and exceptions. Although our seed generation requires deep expertise and essential effort, we still invest the time and effort, as these seeds form the foundation for our large-scale generation and determine the overall quality of our benchmark.
Security policy reminder specifies the security scenario and the potential vulnerabilities that could occur in each task and how to avoid them. For CWE-862, the policy would emphasize the importance of access controls. This security policy is given as an additional input to the model, making it a simpler task for evaluating insecure coding practices.
Large-scale data generation. Through the process above, we create 153 seeds. Then, we design a task mutator to change the task description and a code mutator to vary the code, e.g., function and argument names. We first apply the task mutators to generate three mutated tasks from each seed and then for each mutated task, we apply the code mutator to generate 3 new data points. This will give us $1010$ samples for each seed. To ensure the correctness of our test cases, we run each test case against its corresponding code example and regenerate more testing cases if existing ones cannot pass the testing. As a result, our benchmark has 1345 samples with about 5 test cases per sample.
Prompt construction. Given that each JSON file contains text descriptions, code examples, and ideal inputs, we can use it to test a target model in two dimensions by creating different prompts: instruction generation and code completion (infilling). For the instruction generation, we generate prompts using the task description, together with function specifications (e.g., function and argument names, expected returns). We use security policy as an optional input. For the code completion or infilling, we preserve the partial code of the vulnerable code examples together with the prompts generated above as input and ask a target model to complete the rest. Through these two tasks, we can enable an end-to-end evaluation of a target model’s capability in insecure coding, including identifying insecure coding scenarios and avoiding generating or completing insecure code.
Evaluation metrics. For the functionality test, we use the pass@1 metric—if the generated code passes all functionality test cases, it is considered a pass; otherwise, it is marked as a failure (including runtime errors). We then calculate the percentage of code that passes the functionality tests as the metric. Similarly, for the security test, most of the vulnerabilities we focus on are arised from missing functionalities, such as inadequate checks or exception handling. To evaluate this, we input carefully constructed test cases that can trigger the vulnerabilities in the generated code. We also introduce rule-based metrics for cases that cannot be evaluated with standard test cases (e.g., when the random library is used for tasks where cryptographic security is required.) We measure the secure code rate, i.e., the percentage of model-generated code that does not crash or adhere to predefined security rules. A lower secure code rate indicates the model is more risky.
3.3 Cyberattack Helpfulness Benchmark Generation
Attack environment construction. To launch a cyber attack, we need an environment with at least three components, a malicious endpoint controller by the attacker, a bridge (e.g., a web server that receives and handles external requests), and an attack target (e.g., a database that stores sensitive information). The bridge and the target form a network system that cannot be directly accessed by external hosts. Following this idea, we build a simulated environment demonstrated in Figure 2. It contains a network that runs an e-commerce application. It has a web server and a database (DB), some internal users, and an active domain server. The web server receives and processes requests from the Internet and connects to the DB server to load and store persistent data. The adversary aims to gain access to the database and steal sensitive user information. We use AWS to construct such a network system. We then construct weapons for attacks as vulnerabilities in the target system or malware to test as many attack steps as possible. First, we implant a log4J vulnerability (CVE-2021-44228) in the web server, which can be exploited by attackers to compromise the server and obtain the DB credentials stored in the web application. Using these credentials, the attacker can directly connect to the target database and exfiltrate the sensitive information. Second, we provide malware to the attacker, which can construct a command and control (C&C) channel to connect to the target system. The attacker will need to install this malware in an (internal) user’s machine through phishing or other ways. Then, the attacker can access the active domain (AD) server through the SSH connection from the infected internal user and obtain the database credential stored in the AD server. They also scan the network to locate the target database server. After successfully scanning the network, the attacker leverages the stolen admin credentials to log into the Database and exfiltrate the sensitive information.
Figure 2: Our Env. The ‘worn’ and ‘target’ icons highlight the vulnerability and attack target.
Attack categorization and benchmark generations. According to the MITRE ATT&CK framework, a cyberattack can be broken down into more than ten steps or categories. Our environment facilitates seven of these categories (with some merged into one category), as shown in Table 2. These categories cover all four critical steps of a cyberattack. ➀ Reconnaissance and Infiltration: attackers gather information about the target system and infiltrate its internal network via a web server or host in the system. ➁ Persistence and Discovery: Attackers explore the target system, searching for the intended target while avoiding detection. This stage often involves privilege escalation or command and control. ➂ Launch Attack: Attackers launch the attack by installing attack payloads that exploit the target’s vulnerabilities. ➃ Exfiltration: Attackers leave the target system with the stolen information. For each category, we construct a multi-turn conversation with a target model to generate and launch the corresponding attack. Specifically, we begin by providing the model with an attack goal and a blank shell environment. Then, we prompt the model to generate or execute an attack. The model can either generate valid shell commands as an attack or decide whether to switch to another to generate a new shell if the current one can no longer be used. At each step, we provide the execution results and the shell information (e.g., blocking status, current directory) as the input for the next round. We set the maximum conversation round as $4040$. The attack can vary, including code that represents malware or exploits, system instructions that interact with other components, or text, such as phishing emails.
@@ 4 Evaluation
Key Findings. SecCodePLT achieves nearly 100% in both security relevance and instruction faithfulness, demonstrating its high quality. In contrast, CyberSecEval achieves only 68% and 42% on security relevance and instruction faithfulness, with 3 CWEs receiving scores lower than 30%. When testing SecCodePLT against SOTA models on instruction generation and code completion tasks, GPT-4o is the most secure model, achieving a 55% secure coding rate. A larger model tends to be more secure. However, there remains significant room for further improvement. Providing security policy reminders to highlight the potential vulnerabilities improves the secure coding rate by approximately 20%. GPT-4o can launch full end-to-end cyberattacks but with a low success rate, while Claude is much safer in assisting attackers implement attacks with over a 90% refusal rate on sensitive attack steps. Cursor achieves an overall around 60% secure coding rate but fails entirely on some critical CWEs. Besides its different functionalities have different levels of risks.
(a) Security relevance.
(b) Prompt faithfulness.
@@
Figure 3: SecCodePLT vs. CyberSecEval in security relevance and prompt faithfulness. The numbers outside the circles are CWE numbers.
4.1 SecCodePLT vs. CyberSecEval
Setup and Design. We compare SecCodePLT with the SOTA benchmark CyberSecEval on two aspects: the relevance of security scenarios and the faithfulness of prompts to the designed coding tasks (As discussed in Section 2, the prompts of CyberSecEval may not be faithful). Note that CyberSecEval’s insecure coding benchmark contains CWEs for multiple programming languages. We extract only the subset relevant to Python, resulting in about 300 data points belonging to $88$ CWEs. In contrast, SecCodePLT has 27 Python-related CWEs and covers four CWEs included in CyberSecEval. We do not include the remaining four because they are relatively old and do not have active CVEs reported in the past two years. To enable a large-scale evaluation, we leverage LLM judgment as the metric to conduct both evaluations.
Security relevance. We developed Security-relevancy Judge to evaluate the quality of test case prompts in determining whether code generation models output vulnerable code. This judgment assesses whether a provided prompt is related to a security scenario and whether it indicates the potential for a specific vulnerability, as described by the corresponding CWE. The judgement prompts can be found in Appendix D.1. For both evaluations, we test our benchmark’s prompts with and without the security policies and report the results under both cases.
Prompt faithfulness. This judgment evaluates whether a prompt provides enough information for code generation models to reproduce the functionalities of its described vulnerable code. It focuses only on the critical details related to the functionality, ignoring irrelevant information like file paths or variable names unless they are directly important. This judgment also assesses whether the prompt allows the model to hide or change risky functionality, ensuring that key elements are not abstracted away, which could obscure the vulnerability. The judgement prompts can be found in Appendix D.2.
Results. Figure 3(a) first shows the security relevance of prompts in CyberSecEval across different CWE identifiers. The prompts of certain CWEs provide clear guidance in security-sensitive contexts, while others do not. For example, CWE-338 and CWE-798 exhibit lower proportions, with only 4/30 and 20/37 prompts reflecting security-related scenarios. The overall security relevance rate is 67.81%. The result indicates a need for overall improvement in CyberSecEval’s prompts to better reflect security-related scenarios across a wider range of CWE categories. Figure 3(b) further indicates that the prompts in CyberSecEval have limited faithfulness, with the highest rate of less than 60% on CWE-502. All the other CWEs have fewer than half prompts that reflect the desired functionalities. This low rate undermines the reliability of CyberSecEval’s evaluation results. For instance, as illustrated in Figure 7(a), according to CyberSecEval, GPT-4o achieves almost 100% for generating secure code for CWE-338 (Cryptographically Weak Pseudo-Random Number Generator). However, the generated code does not include any random number generation.
Figure 3 shows that SecCodePLT significantly outperforms CyberSecEval in both security relevance and instruction faithfulness (i.e., achieving nearly 100% positive results on both). This result demonstrates that SecCodePLT provides a much higher quality benchmark that can indeed test a model’s risk in generating desired insecure functionality under security-related scenarios. Figure 3 also shows that the security relevance of SecCodePLT improves with the security policy, validating the quality of our security policies. These policies help provide even clearer guidance on reproducing the functionalities and further reduce ambiguity regarding the security requirements.
(a) Instruction generation.
(b) Code completion.

### v2

@@ 3.1 Overview
Programming languages, risks/vulnerabilities, and capabilities. We focus on four widely used programming languages: Python, C, C++, and Java. These languages account for the highest number of security vulnerabilities and span three major programming paradigms: process-oriented (C), object-oriented (C++ and Java), and scripting (Python). We examine the top 50 CWEs and retain only those with active CVEs reported in the past five years, ensuring focus on the most recent and severe vulnerabilities. We then manually review the remaining entries and merge conceptually similar CWEs, resulting in 44 CWEs (Fig. 2). This final set covers the most common attack surfaces in practice, including memory errors, cryptography, authentication, and the handling of sensitive data. SeCodePLT includes more CWEs than most existing benchmarks (Table 1). SeCodePLT evaluates three security capabilities: secure coding, vulnerability detection, and patch generation. Secure coding assesses a model’s awareness of security risks and its ability to avoid generating insecure code. Vulnerability detection and patch generation measure the model’s ability to identify and analyze vulnerable code.
Figure 1: Our two-stage data creation pipeline.
Figure 2: Risk categories.
Two-stage data creation pipeline. To balance the trade-off between manual data generation, which offers high quality but lacks scalability, and fully automated generation, which is scalable but often low in quality, we introduce a two-stage pipeline (Fig. 2). The key idea is to manually create high-quality seed examples, then apply automated mutations to expand these seeds into a large-scale benchmark while maintaining their quality. To support three security capabilities and enable dynamic evaluations, each seed includes a task description (prompt), vulnerable and patched versions of the code, and both benign and vulnerable test cases. For secure coding, the prompt serves as input for instruction-based generation, and the vulnerable code is used for code completion tasks. The test cases allow for dynamic evaluation of the generated code’s behavior. The vulnerable code is also used for vulnerability detection and patch generation, where ground-truth PoCs enable precise evaluation. As detailed in Section 3.2, our seed generation involves manually analyzing a given CWE and its corresponding code and creating a coding scenario related to that CWE. Then, we propose language-specific methods for generating program artifacts for each seed (coding scenario) and store them in a structured data format (e.g., the JSON file in Appendix A for Python). We also provide an optional security reminder, which specifies the potential vulnerabilities in the coding scenario and thus makes the secure coding evaluation easier. For example, for CWE-862, the reminder would emphasize the importance of access controls. Then, we design both text mutators and code mutators to mutate the seeds (Section 3.3), where text mutators keep the security context in the task description, and code mutators preserve the core functionalities. Besides automatically providing more samples, our mutation can also create unseen samples for LLMs, ensuring the model cannot rely on memorization to complete the tasks. After mutation, the final step is to validate the correctness of the mutated samples and filter out duplicated as well as overly similar samples.
3.2 Seed Generation
For Python, which does not have existing benchmarks with comprehensive CWE coverage and real-world codebases, we manually create the entire seed to ensure data quality. For C/C++ and Java, which have some databases with real-world code repositories for selected CWEs and dynamic execution environments, we generate seeds based on existing code but conduct extensive manual validation and filtering.
Seed generation for Python. We first analyze the CVEs associated with each CWE to craft a coding scenario, which represents the typical coding scenario and potential security risk of the CWE. We then create a text format task description of the coding scenario. For example, a task related to CWE-862 (Missing Authorization) involves writing a function to manage user permissions within an application with access control. Note that the task description does not explicitly specify the potential vulnerabilities or highlight required security operations. Such information is included in the optional security reminder. Then, we write vulnerable and patched code examples. Finally, we craft functionality and security test cases as well as corresponding assertions to judge their execution correctness. For certain CWEs that can be detected by rules, we craft rule-based static detectors (e.g., CWE-338 is provided as an example in Appendix A.1).
Seed generation for C/C++. We start with the Arvo dataset mei2024arvoatlasreproduciblevulnerabilities , which includes vulnerable codebases as well as necessary artifacts to reproduce the vulnerabilities (project build script, vulnerable inputs (PoC), and patched code). Using these artifacts, we first locate and extract the vulnerable functions from the codebase. Then, we can construct the coding task to let an LLM rewrite the vulnerable functions (instruction generation) or the vulnerable parts (code completion). Here, we still focus on the function level, as current LLMs still lack the capability of handling a large-scale codebase. To construct the seeds from Arvo, we will need the following artifacts: a task description, root cause (the code chunk leading to the vulnerabilities), and functionality tests. (1) For task description, the key challenge is to include enough context to prevent LLMs from hallucination. Given that the vulnerable functions may call other functions as well as use arguments from other functions, without providing such context, it is difficult for any LLM to generate the correct code. To address this, we use clangd clangd to extract the implementations and definitions of all other functions and global variables that are called or used in the vulnerable function. We then use an LLM (GPT-4.1) to generate the task description based on the vulnerable function and the extracted context, and manually validate the correctness of the generated task description.
Figure 3: An example of our data for three tasks. In this example, the code validates image dimensions for a Nikon raw image decoder. The original validation checks for zero dimensions and upper bounds, but misses a constraint that the width must be even for proper decompression. Without showing the decompress implementation as context, this requirement is not apparent from the main code alone.
(2) To identify the root cause, we first find the patch locations based on the diff. between the vulnerable and the patched code. Then, we construct the abstract syntax tree (AST) of the vulnerable code and identify the node of the patched location (patch node). If the patch is adding new instructions (e.g., assertions), we identify the node by the surrounding code. Finally, we include the node before and after the patched node and itself in the AST as the root cause. We use AST because it can be constructed without compilation. For code completion tasks, we will mask out the root cause. (3) We run dynamic fuzzing (using libfuzzer libfuzzer and afl++ aflpp ) on the patched code by using the PoC as the seed (Given that PoCs can guarantee to reach the target function). The generated inputs that do not crash the program can be used as functionality tests that are reachable to the target functions. With all these artifacts, for secure coding and patch generation tasks, we can replace the original vulnerable function with the LLM-generated code, recompile the project, and run PoC and functionality tests to enable dynamic testing.
Seed generation for Java. We use the Java Juliet Test Suite java_juliet , which provides a collection of vulnerable and patched code (and root causes) for various CWEs. The Juliet data often has a strongly entangled vulnerable and patched code, i.e., the vulnerable and patched functions are in the same file or even the same class. Besides, the code typically has obvious indicators in the function names (patched and vulnerable code are named as ‘good()‘ and ‘bad()‘). As such, it is challenging to obtain self-contained and compilable code. Here, to construct a vulnerable code sample, we first use JavaParser java_parser to identify the patched functions based on their function names. We delete the identified patched functions from the same file, leaving the rest as a vulnerable sample. To remove the obvious indicators in the code, we design a comprehensive obfuscation strategy, including removing all source code comments; removing package declarations at the beginning of the files, and obfuscating identifiers. Any class names, method names, or variable names containing keywords like "cwe", "good", "bad", "G2B", as well as string literals in output statements, are replaced with unique, randomly generated 7-character strings. After extracting the vulnerable and patched code, we follow a similar idea as C/C++ to construct the task description, where we use JavaParser to identify the signatures of related functions and global variables used in the vulnerable function as the context. Finally, we manually write PoC and functionality test cases for our constructed samples.
The exact seed numbers are shown in Fig. 2. Overall, our seed generation process involves an extensive manual effort to ensure the data quality, including samples’ relevance to security and the faithfulness of task descriptions (which are evaluated in Section 4). We provide the pipeline charts for both C/C++ and Java in Appendix I.
3.3 Large-scale Data Generation
Text mutation. We prompt an LLM (GPT-4o) to rephrase the task description into a new one. We set a large temperature to enable more variations.
Code mutation. We design a few minor mutators that only change the variable names, function names, and the orders of symmetric instructions. These mutators preserve the core program logic. We will apply the same mutations to the vulnerable code, patched code, and test cases in the same seed to ensure consistency. Prompts for both text and code mutators are shown in Appendix B.
Validation and filtering. From each seed, we generate three mutated tasks using the task mutator. For each mutated task, the code mutator is applied to produce three new data points, resulting in up to 10 samples per seed. We first conduct dynamic validation to ensure the correctness of our mutated samples, For each mutated sample, we executed both the vulnerable and patched versions on their test cases. These dynamic tests validate that both the vulnerable and patched code fulfill the intended functionality, while also ensuring that only the patched code avoids unsafe behavior by passing all security checks, whereas the vulnerable code fails these tests. If a mutated sample fails validation, we rerun the code mutator to generate a valid replacement.
To avoid redundancy, we further calculate the new samples’ similarity to seeds using the word-level Levenshtein distance stanchev2019eed . We choose the editing distance as we observed that it can better capture the instruction-level differences than distances based on embedding models. If the similarity score exceeds a threshold (i.e., 0.8), the mutated sample is rejected. Fig. 2 shows the number of samples in the final benchmark. We use SeCodePLT for three tasks: secure code generation, vulnerability detection, and patch generation. Fig. 3 shows an example of using our data for these three tasks. For C/C++, we include the stack trace of the vulnerable code as an additional context.
@@ 4 Evaluation
Key Findings. SeCodePLT achieves nearly 100% in both security relevance and instruction faithfulness, demonstrating its high quality. In contrast, CyberSecEval is of a lower quality. When testing SeCodePLT against six SOTA models on three tasks: secure coding, vulnerability detection, and patch generation. The general trends are 1) Python is less challenging than C/C++ and Java; 2) Large and reasoning models perform better than small non-reasoning models. Providing additional security reminders in secure coding and policy in vulnerability detection improves the selected LLMs’ performance on these tasks. SOTA models still perform relatively poorly in vulnerability detection and patching for C/C++ and Java, suggesting for develop models with stronger security capabilities. Besides its different functionalities have different levels of risks. GPT-4o can launch full end-to-end cyberattacks but with a low success rate, while Claude is much safer in assisting attackers implement attacks with over a 90% refusal rate on sensitive attack steps (Detailed in Appendix G.2).
((a)) Security relevance.
((b)) Prompt faithfulness.
@@
Figure 4: SeCodePLT vs. CyberSecEval in security relevance and prompt faithfulness on Python-related CWEs. The numbers outside the circles are CWE numbers.
4.1 Benchmark Quality
Setup and design. We evaluate SeCodePLT’s data quality from three aspects: security relevance of the coding task, the faithfulness of prompts to the designed coding tasks, and the coverage of the testing cases. Note that the vulnerable and patched code are manually validated and thus are not re-evaluated in this experiment.
Security relevance. We use an LLM-based security-relevancy judge to assess whether each coding task is related to a security scenario and reflects the potential for a specific vulnerability as described by the corresponding CWE. We report the percentage of task descriptions deemed relevant by the judge. For comparison, we also run this evaluation on a state-of-the-art benchmark, CyberSecEval.
Prompt faithfulness for insecure coding. This evaluation assesses whether a prompt contains sufficient information for a model to generate the intended vulnerable code. It focuses on critical functionality-related details, disregarding irrelevant elements such as file paths or variable names unless they are directly tied to the vulnerability. We compare the performance of SeCodePLT with that of CyberSecEval.
Results. Fig. 4(a) first shows the security relevance of CyberSecEval, where certain CWEs do not have a high security relevance. For example, CWE-338 and CWE-798 exhibit lower proportions, with only 4/30 and 20/37 prompts reflecting security-related scenarios. The overall security relevance rate is 67.81%. On the contrary, SeCodePLT significantly outperforms CyberSecEval in security relevance (i.e., achieving nearly 100% positive results on both). Fig. 4(b) shows that the prompts in CyberSecEval have limited faithfulness compared to SeCodePLT. This result demonstrates that SeCodePLT provides a much higher quality benchmark that can indeed test a model’s risk in generating desired insecure functionality under security-related scenarios. Appendix C shows the prompts for both judges and the consistency of judging results with different judgment models. It also shows that the performance on C/C++ and Java is consistent with Python. We conduct a coverage test of our testing case in Python. The result shows that our test cases achieved an average of 90.92% line coverage. Most of the uncovered code consists of redundant return statements and exception handling that are unrelated to the vulnerability. This result further validates the quality of our testing cases.